OpenStudy (anonymous):

A certain hockey player scores on 18% of his shots. What is the probability he will score for the first time on his fifth shot? What is the probability that it will take him fewer than 3 shots to score? How many shots should he expect to take before scoring?

OpenStudy (anonymous):

The probability that Sam will score above a 90 on a mathematics test is 4/5. What is the probability that she will score above a 90 on exactly 3 of the 4 tests this quarter? P(Sam will score >90 on exactly 3 of the 4 tests) = P(Sam fails in 1st and succeeds in others : 2, 3, 4) + P(Sam fails in 2nd and succeeds in others : 1, 3: 4) + P(Sam fails in 3rd and succeeds in others : 1, 2, 4) + P(Sam fails in 4th and succeeds in others : 1, 2, 3) + each term has the same value : P(Sam will score >90 on exactly 3 of the 4 tests) = 4 x (4/5)(1-4/5)^3 = 4² /5^4 = 16/625 = 0.64 2) A certain hockey player scores on 18% of his shots let's pose : a = 0.18 P(he will score for the first time on his fifth shot) = P(he fails on shots 1 to 4) = (1- 0.18)^4 # 0.452 P(it will take him fewer than 3 shots to score) = P( 1st shot ok) + P(1st shot fails and 2nd ok) = a + (1-a)a = 0.18 + 0.82 x 0.18 = 03276 = 32.76% How many shots should he expect to take before scoring? this question is a bit more difficult : P(player fails n times before first success) = a(1-a)^n this introduces the random variable X which is the amount of successive failures before 1st success P(X=n) = a(1 - a)^n and so the definition of the mathematical expectation is E(X) = sigma from n=1 to infinite of ( n a (1 - a)^n ) the calculation (it's a "classical" known calculation) gives E(X) = 1/a # 5.55 hope it'll help!!

OpenStudy (anonymous):

I still don't understand how to do the last part...

OpenStudy (anonymous):

which one are we doing?

OpenStudy (anonymous):

or all?

OpenStudy (anonymous):

The hockey player question, I'm on the last part and unsure how to even start the answer

OpenStudy (anonymous):

the expected number of shots?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if the probability he makes the shot is \(p\) then the expected waiting time is \(\frac{1}{p}\)

OpenStudy (anonymous):

so your answer is \[\frac{1}{.18}\] whatever that is

OpenStudy (anonymous):

why is there a waiting time though?

OpenStudy (anonymous):

and why is it 1 divided by P?

OpenStudy (anonymous):

it is just called "waiting time" is all how many shots do you expect to have to make before you make one

OpenStudy (anonymous):

for example if you roll one die, the probability you get a "2" is \(\frac{1}{6}\) so you expect to have to roll \(6\) times to get a \(2\)

OpenStudy (anonymous):

if you are asking why, to derive the formula requires a summation trick from calculus probably not a able to explain it unless you have had one semester of calc at least

OpenStudy (anonymous):

ok I'm just a little confused because 1 divided by 0.18 is 5.55 but you can't have a decimal point in the number of shots you're taking. Do I just round or...?

OpenStudy (anonymous):

an expected value is a kind of average

OpenStudy (anonymous):

if you have a bunch of numbers, the average of those numbers does not have to be a whole number or even one of the numbers given

OpenStudy (anonymous):

Oh. Okay thanks

OpenStudy (anonymous):

yw